The following function represents a mass undergoing simple harmonic motion.

x(t)=5.0cos(0.40t+0.10) in SI Units

A) Determine the amplitude

B) Determine the angular frequency.

C) Determine the frequency

D) Determine the period

E) Determine the phase constant.

F) Determine the function which represents the velocity as a function of time.
Answer #1

x(t)=5.0cos(0.40t+0.10)

2,5π 7,5m 10π 12,5π 17,5

A) Determine the amplitude

A= 5.0

B) Determine the angular frequency.

w=0.4

C) Determine the frequency

f=w/2pi=0.4/2 pi=0.0636

D) Determine the period

T=1/f=1/0.636=15.707

E) Determine the phase constant.

phase =0.1

F) Determine the function which represents the velocity as a function of time.

V=\frac{\delta X}{\delta t}=-2sin(0.4t+0.1)